Ask
your own question, for FREE!
Mathematics
20 Online
From a standard deck of cards, what is the probability that you draw a diamond, replace it and then draw a club? A) 1 16 B) 1 4 C) 1 2 D) 1
Still Need Help?
Join the QuestionCove community and study together with friends!
A: 1/16
On the first draw, you have a 1/4 chance of drawing a card from 1 of the 4 suits. In other words, you have a 13/52 chance in drawing one of the 13 cards in a suit. 13/52=1/4. If you "replace" the card you just drew (i.e., put it back in the deck), then you'll have a 13/52 chance of drawing another card of a particular suit. Because the two draws are independent, you can multiply the two probabilities, so the probability of getting a diamond and then a club is 13/52*13/52 = 1/4 * 1/4 = 1/16.
oh you posted this loads of time ha sorry
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
13 hours ago
0 Replies
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal